# Improving satellite imagery segmentation using multiple Sentinel-2 revisits

### v1

@@ 2 Related Work
Figure 1: Sample Image-Mask Pair
@@ 3.2 Multi-Temporal Input
As mentioned above, a unique feature of satellite imagery is its multi-temporal aspect. Satellites generate multi-temporal images because they continuously orbit the Earth, re-visiting the same location repeatedly. This results in a unique problem and opportunity for computer vision models using satellite imagery. We identify and evaluate four different general strategies to handle this multi-temporal input in the context of fine-tuning pre-trained models.
@@ 4 Results
We conducted experiments to evaluate the performance of three architectures – SWIN, ViT and U-Net – along with their respective pre-trained weights on the substation segmentation task outlined in Section 3.1. Of particular interest was the question of which of the possible methods for utilizing revisits (Section 3.2) leads to highest performance. We also performed additional experiments to show the effectiveness of pre-training. We then show how our results replicate on the PhilEO building density estimation task (see Section 3.1).
@@ 4.1 SWIN outperforms ViT and U-Net
Model Input Dimension Temporal Aggregation IoU
@@
Table 3: Comparison of models using different strategies for utilizing multi-temporal input on power-substation segmentation task. The input images to the ViT, U-Net and SWIN models had 3, 13 and 9 channels respectively.
@@ 4.2 Incorporating multiple revisits improves performance
Our main question of interest is how to best utilize revisits as a unique feature of satellite imagery. We conducted experiments to evaluate and compare the different techniques of handling multi-temporal input as defined in Section 3.2. Table 3 shows performance for all four revisit techniques for each of the three pre-trained models. Combining the latent intermediate features using a temporal max operation results in a 14%, 17% and 5% lift in performance (over the single-image baseline) for the ViT model, U-Net and SWIN model respectively. Additionally, we find that computing a median image from all the revisits also improves the performance by a few points as compared to using a single image. Interestingly, using each single revisit to augment the dataset doesn’t lead to any noticeable improvement in performance over using just any one single image out of all the revisits; this is in contrast to what was found by Ayala et al. (2022).
@@ 4.3 PhilEO-Downstream Dataset
To test the generality of our claims, we ran the experiments outlined in Sections 4.1 and 4.2 on the PhilEO Downstream dataset and show results in table 4. Consistent with our earlier findings, the SWIN model demonstrated superior performance compared to the U-Net and ViT models on the building density estimation task. Once again, the ViT model ranked worst amongst the three models. Also aligned with our earlier results, the latent temporal fusion strategy proved to be the most effective strategy of incorporating multi-temporal images, resulting in a 5%, 12% and 17% improvement in the ViT, U-Net, and SWIN model’s performance respectively compared to using a single image as input. Similarly, using median images led to some improvement over single image inputs. Finally, our best models outperform models from Fibaek et al. (2024), further validating the effectiveness of the temporal fusion strategy (though note the comparison is not exact because we are using a restricted dataset of only locations that contained at least 3 revisits). In total, we find that combining temporal data in the latent space is a universally effective way to utilize revisits in satellite imagery.
@@ 5 Conclusion
Importantly, we took an application-driven approach to this problem by focusing on a dataset and task that is of interest the energy sector, particularly in planning the energy transition (Rolnick et al., 2024). The problems faced in this application are common for many use-cases for remote sensing and geospatial data: a specific task that is not necessarily represented in many existing benchmarks, a small- to medium-sized dataset, and the desire to minimize compute costs. These constraints create a setting where fine-tuning pre-trained models is a natural solution. By identifying the best way to utilize revisits in the context of fine-tuning specific pre-trained remote sensing models, we believe we are providing valuable insights for many applied researchers.
We also believe our work should drive future studies exploring the most effective ways of fusing representations in the latent space, as we only explored a small subset of possible ways here. Another interesting question is that of how many revisits are necessary, and does performance plateau after a certain number of revisits (when using revisits as data augmentation, Ayala et al. (2022) saw performance increase as a function of revisits up to 3). Finally, we expect our findings to apply to tasks involving static features such as buildings or roads. However it is possible that revisits taken within a short period of time (e.g., the multiple images Sentinel-2 takes in the course of a month) may help even in the context of more temporally varying targets, such as vegetation classification.
@@
Appendix A Normalization Techniques
We experimented with three different data preprocessing techniques for scaling pixel values of input images: standardization, normalization and scaling down by a constant.
@@ Appendix A Normalization Techniques
For standardization, we used a large random sample of the dataset to compute the mean and standard deviation. Since normalization is sensitive to outliers, we used the 1st and 99th percentiles in place of minimum and maximum values. Scaling down by a constant, which in this case is similar to normalization since our input image data didn’t have negative values, worked best with a constant value of 4000. Different strategies worked well for different pretrained models. For the SWIN model and the U-Net, scaling with a constant yielded the best results, while for the ViT model, standardization performed the best. This was possibly because the ViT encoder was pretrained on ImageNet where z-score standardization is applied on input data.
@@ Appendix C Incorporating more channels boosts performance
Table 7: Comparing different ViT encoders on the power-substation segmentation task. Models were trained on single timepoint that was scaled with a constant(Appendix A)
@@ Appendix E ViT Decoder Architecture
The ViT Encoder was applied in parallel to the multiple timepoints following which the individual embeddings were fused using a temporal max operation. These were passed to the decoder(Figure 5). The decoder consisted of sequential up-scaling blocks that transformed the feature embedding from the encoder into an output mask of shape 224x224. Each up-scaling block was composed of ConvTranspose, Conv2d, BatchNorm, Dropout and ReLU layers.

### v2

@@ 2 Related Work
Figure 1: Images from multiple revisits along with mask
@@ 3.2 Multi-Temporal Input
As mentioned above, a unique feature of satellite imagery is its multi-temporal aspect. Satellites generate multi-temporal images because they continuously orbit the Earth, re-visiting the same location repeatedly. This results in a unique problem and opportunity for computer vision models using satellite imagery. We identify and evaluate five different general strategies to handle this multi-temporal input in the context of fine-tuning pre-trained models.
@@
  • Output Fusing of Multi-Image Input: Generate individual pre-threshold segmentation maps for each revisit, then threshold the median of these images. This method also utilizes all revisits for each training step, but does not require alterations to the internal model layers and is therefore more agnostic to model architecture.
@@ 4 Results
We conducted experiments to evaluate the performance of three architectures – SWIN, ViT and U-Net – along with their respective pre-trained weights on the substation segmentation task outlined in Section 3.1. Of particular interest was the question of which of the possible methods for utilizing revisits (Section 3.2) leads to highest performance. We also performed additional experiments to show the effectiveness of pre-training. We then show how our results replicate on the PhilEO building density estimation task (see Section 3.1). For all the experiments, we report mean and standard error over 5 runs with random seeds.
@@ 4.1 SWIN outperforms ViT and U-Net
Model Temporal Dimension Temporal Aggregation IoU
@@
Table 3: Comparison of different strategies for utilizing multi-temporal input on power-substation segmentation task. All the models used a pretrained encoder and respective multi-spectral inputs (Appendix B).
@@ 4.2 Incorporating multiple revisits improves performance
Our main question of interest is how to best utilize revisits as a unique feature of satellite imagery. We conducted experiments to evaluate and compare the different techniques of handling multi-temporal input as defined in Section 3.2. Table 3 shows performance for all five revisit techniques for each of the three pre-trained models. Combining the latent intermediate features using a temporal max operation results in a 14%, 17% and 5% lift in performance (over the single-image baseline) for the ViT model, U-Net and SWIN model respectively. Additionally, we find that computing a median image from all the revisits also improves the performance by a few points as compared to using a single image. Interestingly, using each single revisit to augment the dataset doesn’t lead to any noticeable improvement in performance over using just any one single image out of all the revisits; this is in contrast to what was found by Ayala et al. (2022).
@@ 4.3 PhilEO-Downstream Dataset
To test the generality of our claims, we ran the experiments outlined in Sections 4.1 and 4.2 on the PhilEO Downstream dataset and show results in table 4. Consistent with our earlier findings, the SWIN model demonstrated superior performance compared to the U-Net and ViT models on the building density estimation task. Once again, the ViT model ranked worst amongst the three models. Also aligned with our earlier results, the latent temporal fusion strategy proved to be the most effective strategy of incorporating multi-temporal images, resulting in a 5%, 12% and 17% improvement in the ViT, U-Net, and SWIN model’s performance respectively compared to using a single image as input. Similarly, using median images led to some improvement over single image inputs. As with our substation results, the output fusion method performed inconsistently, occasionally performing even worse than the single image baseline (though note we did experience optimization issues with this approach, especially when using the U-net architecture). Finally, our best models outperform models from Fibaek et al. (2024), further validating the effectiveness of the temporal fusion strategy (though note the comparison is not exact because we are using a restricted dataset of only locations that contained at least 3 revisits). In total, we find that combining temporal data in the latent space is a universally effective and robust way to utilize revisits in satellite imagery.
@@ 5 Conclusion
Importantly, we took an application-driven approach to this problem by focusing on a dataset and task that is of interest to the energy sector, particularly in planning the energy transition (Rolnick et al., 2024). The problems faced in this application are common for many use-cases for remote sensing and geospatial data: a specific task that is not necessarily represented in many existing benchmarks, a small- to medium-sized dataset, and the desire to minimize compute costs. These constraints create a setting where fine-tuning pre-trained models is a natural solution. By identifying the best way to utilize revisits in the context of fine-tuning specific pre-trained remote sensing models, we believe we are providing valuable insights for many applied researchers.
We also believe our work should drive future studies exploring the most effective ways of fusing representations in the latent space, as we only explored a small subset of possible ways here. Another interesting question is that of how many revisits are necessary, and does performance plateau after a certain number of revisits (when using revisits as data augmentation, Ayala et al. (2022) saw performance increase as a function of revisits up to 3). Finally, we expect our findings to apply to tasks involving static features such as buildings or roads. However it is possible that revisits taken within a short period of time (e.g., the multiple images Sentinel-2 takes in the course of a month) may help even in the context of more temporally varying targets.
@@
Model Normalization Technique IoU
SWIN Normalization 0.572 $\pm$ 0.0033
SWIN Scaling with a constant 0.582 $\pm$ 0.0026
Table 5: Comparison of different data pre-processing techniques for the SWIN model on power-substation segmentation task. Temporal max operation is used to latently fuse features of each timepoint.
Model Normalization Technique IoU
ViT Scaling with a constant 0.335 $\pm$ 0.0001
ViT Standardization 0.382 $\pm$ 0.0001
Table 6: Comparison of different data pre-processing techniques for the ViT based model on power-substation segmentation task. Single image input is used for this experiment.
Appendix A Data Pre-processing Techniques
We experimented with three different data pre-processing techniques for scaling pixel values of input images: standardization, normalization and scaling down by a constant.
@@ Appendix A Data Pre-processing Techniques
For standardization, we used a large random sample of the dataset to compute the mean and standard deviation. Since normalization is sensitive to outliers, we used the 1st and 99th percentiles in place of minimum and maximum values. Scaling down by a constant, which in this case is similar to normalization since our input image data didn’t have negative values, worked best with a constant value of 4000. Different strategies worked well for different pretrained models. For the SWIN model and the U-Net, scaling with a constant yielded the best results(see SWIN results in table 5), while for the ViT model, standardization performed the best(see table 6). This was possibly because the ViT encoder was pretrained on ImageNet where z-score standardization is applied on input data.
@@ Appendix C Incorporating more channels boosts performance
Table 9: Comparing different ViT encoders on the power-substation segmentation task. Single image input was used and images were scaled with a constant(Appendix A)
@@ Appendix E ViT Decoder Architecture
The ViT Encoder was applied in parallel to each timepoint and the individual embeddings were fused using a temporal max operation. The resulting embedding was passed to the decoder(Figure 5) which consisted of sequential up-scaling blocks that transformed the feature embedding into an output mask of shape 224x224. Each up-scaling block was composed of ConvTranspose, Conv2d, BatchNorm, Dropout and ReLU layers.